# MetaAligner: Towards Generalizable Multi-Objective Alignment of Language Models

### v2

@@ Abstract
Recent advancements in large language models (LLMs) aim to tackle heterogeneous human expectations and values via multi-objective preference alignment. However, existing methods are parameter-adherent to the policy model, leading to two key limitations: (1) the high-cost repetition of their alignment algorithms for each new target model; (2) they cannot expand to unseen objectives due to their static alignment objectives. In this work, we propose Meta-Objective Aligner (MetaAligner), a model that performs conditional weak-to-strong correction for weak responses to approach strong responses. MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning. Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 10 state-of-the-art policy models, and outperforms previous alignment methods with down to 15.71$\times$ less GPU training hours. The model also effectively aligns unseen objectives, marking the first step towards generalizable multi-objective preference alignment. This project is open-sourced at: https://github.com/SteveKGYang/MetaAligner
@@ 1 Introduction
The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values. At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) [39, 23, 28], where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model [23].
However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences [4, 25]. These drawbacks motivate further exploration into multi-objective alignment algorithms. Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) [27, 20, 25]. Due to its substantial computational cost [20, 25] and the unstable nature of the PPO [26, 17, 24], other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) [38, 11] or supervised fine-tuning (SFT)-based methods [35, 11], which customized prompting strategies to incorporate multiple reward values into queries explicitly.
The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model. This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models [2, 31, 7, 30]; (2) all target models are statically aligned on pre-determined (e.g. "Helpful", "Harmless", "Honest" [38, 11]) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives. This ignorance leads to poor generalizability of existing multi-objective alignment methods.
Algorithm Paradigm Reward-Free Align. Multi-Objective Align. Policy-Agnostic Align. Generalizability
@@
Table 1: A checklist for key characteristics of previous preference alignment methods and MetaAligner. "Align." denotes "Alignment".
@@
In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access [13]. Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives. The enhanced datasets supervise MetaAligner via conditional weak-to-strong correction. That is, MetaAligner is trained to correct the weak responses to approach the strong responses considering the corresponding alignment objectives, which are incorporated via prompt engineering. During inference, MetaAligner is stacked onto any policy models to perform objective-aware corrections. As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment. The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning [16]. The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives. A checklist of MetaAligner’s advantages over previous methods is presented in Table 1.
@@
Through this work, we make three main contributions: (1) we propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment. It performs multi-objective alignment efficiently, without tuning the policy models or accessing their parameters. Experimental results show that MetaAligner outperforms previous alignment methods with down to 15.71× less GPU training hours and higher stability; (2) we utilize MetaAligner to exert zero-shot preference alignment for unseen objectives. To our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment. Experimental results show that MetaAligner can simultaneously perform effective alignment for 6 unseen objectives while maintaining performance on aligned objectives; (3) We examine MetaAligner on 3 preference alignment datasets. Experimental results show that MetaAligner improves win rates on multiple objectives across 10 policy models, substantially enhancing responses of state-of-the-art foundation models such as GPT-3.5 [22] and Claude-3 [3].
@@
Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF [20], (B) SFT-based methods [35, 11], and (C) Meta-Objective Aligner. SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
@@ 2 Multi-Objective Alignment of Language Models
In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplays such as contradiction (e.g. "Helpful" and "Harmless") and dependence (e.g. "Correct" and "Informative"). Multi-objective preference alignment tackles this challenge by optimizing multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$. $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry [6] model of preferences. Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
$P_{\mathcal{D}}(y_{1}\succ y_{2}|q,i)=\frac{1}{1+exp(-(r_{i}(q,y_{1})-r_{i}(q,y_{2})))}$ (1)
where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards [27, 20] with a KL-divergence regularization (see Figure 1 (A)):
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector. Another paradigm [35, 11] directly built alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[P_{\pi_{\phi}}(y|q,\mathbf{R}(q,y))\right]$ (3)
where objectives and their corresponding reward values are described with text markers and combined into queries with a static prompting template. Compared to MORLHF, SFT-based multi-objective alignment is proven more cost-efficient and training-stable.
@@ 3 Meta-Objective Aligner
This section introduces the proposed methodology for MetaAligner, which follows a 3-stage paradigm: (1) dynamic multi-objective dataset construction; (2) a 3-step model training process; (3) generalizable inference for alignment. The paradigm is illustrated in Figure 1 (C).
3.1 Dynamic Multi-Objective Dataset
@@
where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text descriptions for the target objectives, and $t\in\{equal,better\}$ depends on the building subset. The detailed construction process is described in Algorithm 1. A full list of the utilized prompting templates is presented in Appendix B.
Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping [35] process in previous SFT-based multi-objective alignment methods.
3.2 MetaAligner Derivation
Based on the dynamic multi-objective dataset, we achieve multi-objective alignment by introducing MetaAligner, a standard conditional seq-to-seq model on top of the original policy model $\pi_{\phi}$, which re-distributes the initial outputs as follows:
@@ 3.2 MetaAligner Derivation
where $\delta_{\theta}$ denotes the MetaAligner module parameterized by $\theta$. We achieve alignment by directly employing SFT on the target response. We have the standard cross-entropy loss as the training objective:
@@
$\mathop{argmin}\limits_{\theta,\phi}$ $\mathcal{L}(\theta,\phi;\mathcal{D})=-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\pi^{*}(y|q)\right]$ (5)
=-= - $\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]-$
$\mathbb{E}_{q\sim\mathcal{D}}\left[log\pi_{\phi}(y_{0}|q)\right]$
We fuse the parameters of the policy model, thus excluding $\phi$ from the weight update process. As only $\theta$ derivatives are calculated during back-propagation, we eliminate the second term in Eqn 5 and simplify the training objective as:
@@
$\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[log\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
@@
The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as samples from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs [2, 22, 3].
3.3 Model Training
In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process. We propose a 3-step paradigm based on the objective function in Eqn. 6:
Step 1: warm-up.
The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies [12, 13] to familiarize the model with identity mapping and facilitate convergence. Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
Step 2: equal-preference modeling.

### v3

@@ Abstract
Recent advancements in large language models (LLMs) focus on aligning to heterogeneous human expectations and values via multi-objective preference alignment. However, existing methods are dependent on the policy model parameters, which require high-cost repetition of their alignment algorithms for each new policy model, and they cannot expand to unseen objectives due to their static alignment objectives. In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner models multi-objective alignment into three stages: (1) dynamic objectives reformulation algorithm reorganizes traditional alignment datasets to supervise the model on performing flexible alignment across different objectives; (2) conditional weak-to-strong correction paradigm aligns the weak outputs of fixed policy models to approach strong outputs with higher preferences in the corresponding alignment objectives, enabling plug-and-play inferences on any policy models, which significantly reduces training costs and facilitates alignment on close-source policy models; (3) generalizable inference method flexibly adjusts target objectives by updating their text descriptions in the prompts, facilitating generalizable alignment to unseen objectives. Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 10 state-of-the-art policy models, and saves up to 93.63% of GPU training hours compared to previous alignment methods. The model also effectively aligns unseen objectives, marking the first step towards generalizable multi-objective preference alignment. This project is open-sourced here.
@@ 1 Introduction
The recent advancements in large language models (LLMs) have focused on generating high-quality responses that align with human expectations and values. At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) [41, 23, 28], where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning (RL) on the target model [23].
However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences [4, 25]. These drawbacks motivate further exploration into multi-objective alignment algorithms. Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) [27, 20, 25]. Due to its substantial computational cost [20, 25] and the unstable nature of the proximal policy optimization (PPO) [26, 16, 24] algorithm, other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) [40, 11] or supervised fine-tuning (SFT)-based methods [36, 11], which customized prompting strategies to incorporate multiple reward values into queries explicitly.
The above methods for multi-objective alignment bear one commonality: the dependence on the policy model’s parameters. This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced policy model, which is incompatible with the increasing sizes and fast iteration of current foundation models [2, 31, 7, 30]; (2) all target models are statically aligned on pre-determined (e.g. "Helpful", "Harmless", "Honest" [40, 11]) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives. This ignorance leads to poor generalizability of existing multi-objective alignment methods.
Table 1: Comparisons between previous alignment methods and MetaAligner on different features. "Policy-Agnostic Alignment" means the alignment algorithm is independent of the target policy model parameters, and "Generalizability" denotes zero-shot alignment capability on unseen objectives.
Algorithm Paradigm Multi-Objective Alignment Policy-Agnostic Alignment Generalizability
@@
In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner models multi-objective alignment into three stages: (1) the dynamic objectives reformulation algorithm reorganizes traditional alignment datasets into dynamic-objective alignment datasets, training MetaAligner to perform flexible alignment across different objectives. It achieves this by incorporating and combining text descriptions of various alignment objectives in a prompt-based manner; (2) the conditional weak-to-strong correction paradigm aligns the weak outputs of policy models to approach strong outputs with higher preferences in the corresponding alignment objectives. During training, MetaAligner is stacked onto policy models to perform objective-aware corrections, where parameters of the policy model are fixed and MetaAligner is optimized with an SFT-based three-step training process: warming up, equal-preference alignment, and contrastive-preference alignment. This paradigm enables MetaAligner to perform plug-and-play inferences on any policy models even without access to their parameters, which significantly reduces training costs and facilitates alignment on close-source LLMs; (3) the generalizable inference method flexibly adjusts target objectives by updating their text descriptions in the prompts. This method can also adapt MetaAligner to unseen objectives and achieve new alignment strategies via in-context learning [15], a new feature with rare previous exploration in alignment of language models. The number of aligned objectives also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives. Table 1 compares key features between MetaAligner and previous methods. As shown, conditional weak-to-strong correction of MetaAligner extends Aligner [13] to multi-objective alignment scenarios, which are not directly solvable by Aligner itself. MetaAligner is also the first multi-objective alignment method to achieve policy-agnostic alignment and generalization to unseen objectives, two key advantages over previous methods such as MORLHF, MODPO, and SFT-based methods.
@@
In summary, our main contributions are: (1) we propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment. It performs multi-objective alignment efficiently, without tuning the policy models or accessing their parameters. Experimental results show that MetaAligner outperforms previous alignment methods and saves up to 93.63% of GPU training hours; (2) we utilize MetaAligner to exert zero-shot preference alignment for unseen objectives. To our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment. Experimental results show that MetaAligner can simultaneously perform effective alignment for six unseen objectives while maintaining performance on aligned objectives; (3) We examine MetaAligner on three preference alignment datasets. Experimental results show that MetaAligner improves win rates on multiple objectives across 10 policy models, substantially enhancing responses of state-of-the-art foundation models such as GPT-3.5-Turbo [22] and Claude-3 [3].
@@
Figure 1: Illustrations of Meta-Objective Aligner, which follows a three-stage paradigm.
@@ 2 Multi-Objective Alignment of Language Models
In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplays such as contradiction (e.g. "Helpful" and "Harmless") [11, 36] and dependence (e.g. "Correct" and "Informative") [35]. Multi-objective preference alignment tackles this challenge by aiming to optimize multiple objectives simultaneously. For each query-response pair, the reward vector is formalized as: $\mathbf{R}(q,y)=[r_{1}(q,y),...,r_{N}(q,y)]^{\mathrm{T}}$, where $q,y$ denote a query and a corresponding response, $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry [6] model of preferences. Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
$P_{\mathcal{D}}(y_{1}\succ y_{2}|q,i)=\sigma(r_{i}(q,y_{1})-r_{i}(q,y_{2}))$ (1)
where $\sigma$ denotes the logistic function and $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards [27, 20] with a KL-divergence regularization:
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}(y|q)}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)\right]-\beta\mathbb{D}_{KL}\left[\pi_{\phi}(y|q)\|\pi_{ref}(y|q)\right]$ (2)
where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$, $\pi_{ref}$ denotes the reference policy model, $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the pre-determined heuristic target preference vector. Another paradigm directly built alignment between multiple reward values and their corresponding response by minimizing an SFT loss for the policy model:
$\mathop{argmin}\limits_{\pi_{\phi}}-\mathbb{E}_{(q,y)\sim\mathcal{D}}\left[\mathop{log}\pi_{\phi}(y|q,\mathbf{R}(q,y))\right]$ (3)
where objectives and their corresponding reward values are described with text markers and combined into queries with a static prompting template. Compared to MORLHF, SFT-based multi-objective alignment is proven more cost-efficient and training-stable [36, 11].
@@ 3 Meta-Objective Aligner
Existing methods for multi-objective alignment generally face challenges in increasing training costs with new policy models and generalization to unseen objectives. To tackle these challenges, we introduce MetaAligner, which follows a three-stage paradigm: (1) dynamic objectives reformulation for building dynamic multi-objective datasets; (2) conditional weak-to-strong correction for model training; (3) generalizable inference for multi-objective alignment. The paradigm is illustrated in Figure 1.
3.1 Dynamic Objectives Reformulation
We propose a dynamic objectives reformulation algorithm to construct a dynamic multi-objective dataset, which triggers MetaAligner’s ability for flexible adjustment of alignment objectives. Specifically, any typical multi-objective preference alignment dataset $\mathcal{D}_{m}$ with $m$ samples and $N$ objectives can be re-organized as $\{q_{i},y_{i1},y_{i2},P_{i}\}_{i=1}^{m}$, where $P_{i}=[p_{i1},...,p_{iN}]^{\mathrm{T}}$ and $p_{ij}\in\{\succ,\prec,\equiv\}$ indicates the preference on $j$-th objective.
Algorithm 1 Dynamic objectives reformulation.
[…]
2:Contrastive subset $\mathcal{D}_{c}$; Equal subset $\mathcal{D}_{e}$.
3:$\mathcal{D}_{c}\leftarrow\varnothing$, $\mathcal{D}_{e}\leftarrow\varnothing$triangleright Initialize the 2 subsets.
4:for $i\in\{1,...,m\}$ do $\triangleright$ Loop on instances.
5:     $\mathcal{O}_{\succ}\leftarrow\varnothing$, $\mathcal{O}_{\prec}\leftarrow\varnothing$, $\mathcal{O}_{\equiv}\leftarrow\varnothing$
6:     for $j\in\{1,...,N\}$ do
7:         if $p_{ij}$ is $\succ$ then $\triangleright$ Collect the objectives where $y_{i1}$ outperforms $y_{i2}$.
8:              $\mathcal{O}_{\succ}\leftarrow\mathcal{O}_{\succ}\cup\{\left<d_{j}\right>\}$
9:         else if $p_{ij}$ is $\prec$ $\triangleright$ Collect the objectives where $y_{i2}$ outperforms $y_{i1}$.
10:              $\mathcal{O}_{\prec}\leftarrow\mathcal{O}_{\prec}\cup\{\left<d_{j}\right>\}$
11:         $\triangleright$ Collect the objectives where $y_{1}$ and $y_{2}$ performs equally.
12:              $\mathcal{O}_{\equiv}\leftarrow\mathcal{O}_{\equiv}\cup\{\left<d_{j}\right>\}$
13:         end if
14:     end for
15:     if $\mathcal{O}_{\succ}\neq\varnothing$ $\triangleright$ Build the training pairs where $y_{i1}$ is used as the target.
16:         $t\leftarrow better$
17:         $\mathcal{O}_{\succ}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\succ})$
18:         $\mathcal{D}_{c}\leftarrow\mathcal{D}_{c}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\succ},t),y_{i1})\}$
19:     end if
20:     if $\mathcal{O}_{\prec}\neq\varnothing$ $\triangleright$ Build the training pairs where $y_{i2}$ is used as the target.
21:         $t\leftarrow better$
22:         $\mathcal{O}_{\prec}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\prec})$
23:         $\mathcal{D}_{c}\leftarrow\mathcal{D}_{c}\cup\{(\mathcal{T}(q_{i},y_{i1},\mathcal{O}_{\prec},t),y_{i2})\}$
24:     end if
25:     if $\mathcal{O}_{\equiv}\neq\varnothing$ $\triangleright$ Build equally-preferred training pairs.
26:         $t\leftarrow equal$
27:         $\mathcal{O}_{\equiv}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\equiv})$
28:         $\mathcal{D}_{e}\leftarrow\mathcal{D}_{e}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\equiv},t),y_{i1})\}$
29:     end if
30:end for
@@ 3.1 Dynamic Objectives Reformulation
where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the concatenation of text descriptions for the target objectives, and $t\in\{equal,better\}$ depends on the current building subset. Details of the dynamic objectives reformulation algorithm are described in Algorithm 1. For an instance within the processing loop (line 2): $\{q,y_{1},y_{2},P\}$, the algorithm performs a two-step reformulation: (1) collect the sets $\mathcal{O}_{\succ}$, $\mathcal{O}_{\prec}$, $\mathcal{O}_{\equiv}$ that includes objectives where $y_{1}$ outperforms $y_{2}$, $y_{2}$ outperforms $y_{1}$, and both perform equally (lines 3-12); (2) for each objective set $\mathcal{O}$, we randomly shuffle the objectives to further trigger the model’s flexible alignment ability, and build query-response pairs based on the corresponding prompting template $\mathcal{T}$ (line 13-27). All prompting templates and examples of the algorithm are presented in Appendix C.
Training on dynamic multi-objective datasets provides three key advantages: (1) instance-level alternation of the target objectives during training enables MetaAligner to perform flexible alignment under different combinations of objectives; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors. (3) the reward-free alignment method (no explicit preference values required) avoids complicated preference-to-reward mapping [36] process in previous SFT-based multi-objective alignment methods.
3.2 Conditional Weak-to-Strong Correction
Based on the dynamic multi-objective training datasets, we train MetaAligner in a conditional weak-to-strong correction manner, which follows an SFT-based training objective and a three-step training paradigm.
3.2.1 Training Objective Derivation
MetaAligner is a standard conditional seq-to-seq model on top of the original policy model $\pi_{\phi}$, which re-distributes the policy model output $y_{0}$ considering objectives $\mathcal{O}$ as follows:
@@ 3.2.1 Training Objective Derivation
where $\delta_{\theta}$ denotes the MetaAligner module parameterized by $\theta$, $t$ depends on the training dataset. Conditional weak-to-strong correction directly trains MetaAligner to align the weak policy model output $y_{0}$ to the strong target output $y$, which has higher preference values in corresponding objectives $\mathcal{O}$. We have the standard cross-entropy loss as the training objective:
@@
$\mathop{argmin}\limits_{\theta,\phi}\mathcal{L}(\theta,\phi;\mathcal{D})$ $=-\mathbb{E}_{(q,y,\mathcal{O})\sim\mathcal{D}}\left[\mathop{log}\pi^{*}(y|q)\right]$ (5)
$=-\mathbb{E}_{(q,y,\mathcal{O})\sim\mathcal{D}}\left[\mathop{log}\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]-\mathbb{E}_{q\sim\mathcal{D}}\left[\mathop{log}\pi_{\phi}(y_{0}|q)\right]$
We fix the parameters of the policy model, thus excluding $\phi$ from the weight update process. In practice, we use the dynamic multi-objective dataset for supervision, where the weak response in each query-response pair is directly leveraged as samples $y_{0}$ from unknown policy models. Therefore, we eliminate the second term in Eqn. 5 and simplify the training objective as:
@@
$\mathop{argmin}\limits_{\theta}-\mathbb{E}_{(q,y_{0},y,\mathcal{O})\sim\mathcal{D}}\left[\mathop{log}\delta_{\theta}(y|\mathcal{T}(q,y_{0},\mathcal{O},t))\right]$ (6)
@@
The above action poses two advantages: (1) the computation resources required for MetaAligner training is detached from policy model size, which enables policy-agnostic and cost-efficient alignment for large policy models; (2) MetaAligner works only via outputs from the policy models, which allows training and inference for alignment on close-source policy models [2, 22, 3].
3.2.2 Three-Step Model Training